OpenStudy (anonymous):

If f(x) = 5x - 9, find the value of f(-2). A. 8 B. -19 C. 23 D. -23

OpenStudy (mathstudent55):

Just evaluate the function after replacing x with -2 f(x) = 5x - 9 f(-2) = 5*(-2) - 9 = ?

OpenStudy (anonymous):

-19 :) B

OpenStudy (mathstudent55):

correct